# How does Alignment Enhance LLMs’ Multilingual Capabilities? A Language Neurons Perspective

### v1

@@ Abstract
Multilingual Alignment is an effective and representative paradigm to enhance LLMs’ multilingual capabilities, which transfers the capabilities from the high-resource languages to the low-resource languages. Meanwhile, some researches on language-specific neurons reveal that there are language-specific neurons that are selectively activated in LLMs when processing different languages. This provides a new perspective to analyze and understand LLMs’ mechanisms more specifically in multilingual scenarios. In this work, we propose a new finer-grained neuron identification algorithm, which detects language neurons (including language-specific neurons and language-related neurons) and language-agnostic neurons. Furthermore, based on the distributional characteristics of different types of neurons, we divide the LLMs’ internal process for multilingual inference into four parts: (1) multilingual understanding, (2) shared semantic space reasoning, (3) multilingual output space transformation, and (4) vocabulary space outputting. Additionally, we systematically analyze the models before and after alignment with a focus on different types of neurons. We also analyze the phenomenon of “Spontaneous Multilingual Alignment”. Overall, our work conducts a comprehensive investigation based on different types of neurons, providing empirical results and valuable insights for better understanding multilingual alignment and multilingual capabilities of LLMs. [^1] The project will be available at https://github.com/NJUNLP/Language-Neurons-Alignment.
1 Introduction
@@
By training on the extensive corpus, large language models (LLMs) demonstrate outstanding language capabilities (Grattafiori et al., 2024; Yang et al., 2024; Liu et al., 2024; Zhang et al., 2025). However, due to the unbalanced pretraining corpus across different languages, LLMs have very uneven performance on high-resource languages and low-resource languages (Huang et al., 2023; Zhu et al., 2023; Zhang et al., 2024). Therefore, researchers have conducted comprehensive explorations to further enhance the multilingual performance of LLMs. A straightforward approach is increasing the proportion of non-English texts during pretraining (Ni et al., 2021; Yang et al., 2024) or performing continual pretraining with multilingual texts (Liu et al., 2021; Ji et al., 2024). But these approaches often entail high computational costs and substantial amounts of multilingual data.
Considering LLMs’ great performance on the high-resource languages, multilingual alignment has emerged as a representative paradigm for enhancing multilingual capabilities by transferring knowledge from high-resource to low-resource languages (Zhao et al., 2024a; She et al., 2024). A representative example is MAPO She et al. (2024), which improves multilingual alignment by utilizing a well-trained multilingual translation model to compute alignment scores based on the conditional generation probability of translating non-English responses into English.
@@
Figure 1: A neuron’s activation probability across different languages. This neuron that exhibits high activation probabilities across multiple but not all languages can’t be correctly categorized under the existing methodology.
Many studies conduct systematic mechanism analyses of the multilingual alignment and LLMs’ multilingual capabilities. Zhao et al. (2024b) split the multilingual processing workflow into three parts: multilingual understanding, resolving tasks, and generating outputs in the target language. This three-stage inference workflow clearly demonstrates how LLMs leverage English as a pivot language to handle multilingualism using a unified pattern. Inspired by the neurobiological underpinnings of human language faculties, Tang et al. (2024) conducts a fine-grained identification by detecting the language-specific neurons. Their results indicate that LLMs predominantly utilize a small subset of neurons to process the particular target language. Furthermore, these language-specific neurons are primarily situated in the model’s top and bottom layers (Tang et al., 2024), which is consistent with the three-stage multilingual workflow of Zhao et al. (2024b). However, we notice a key limitation in the existing language-specific neuron identification methodology: Some neurons are shared across multiple languages but are not entirely language-agnostic. Such neurons are incorrectly categorized as either language-specific or language-agnostic neurons under the existing framework. We present a case study in Figure 1. Furthermore, we aim to systematically explore an important question: Can we better analyze and understand how multilingual alignment enhances the LLMs’ multilingual capabilities from the perspective of language neurons?
In this work, we comprehensively investigate the multilingual alignment of LLMs from the perspective of language neurons, where we deploy MAPO as a representative multilingual alignment algorithm. Considering the above limitation of the existing language-specific neuron identification methodology, we define language neurons as the union of language-specific neurons and language-related neurons, as opposed to language-agnostic neurons. We separately distinguish language-related neurons from both language-specific and language-agnostic neurons, which allows for more precise analyses. Furthermore, we propose a new language neuron identification algorithm in our work, which is able to identify language-related neurons that are in fact shared across multiple languages.
@@
Then we analyze the models before and after alignment, focusing on the changes in different types of neurons. Based on the distributional characteristics of these neurons, we divide LLMs’ internal process for multilingual inference into four parts: (1) multilingual understanding, (2) shared semantic space reasoning, (3) multilingual output space transformation, and (4) vocabulary space outputting. Our findings reveal that different parts exhibit distinct dependencies on different types of neurons, and that multilingual alignment significantly enhances the activation of the corresponding types of neurons across the relevant layers. Additionally, we analyze the important “Spontaneous Multilingual Alignment” (Zhang et al., 2024) phenomenon in LLMs, providing insights into the roles of language-agnostic neurons and language-related neurons shared across languages. For further analysis, we also provide observations about the uniqueness of English and the neuron distributions. Overall, based on different types of neurons, we present empirical results and valuable insights that contribute to a deeper understanding of multilingual alignment and the multilingual capabilities of LLMs.
2 Related Work
2.1 Multilingual Alignment
@@
Conducting pretraining or continual pretraining on the multilingual corpus is a straightforward and effective method to enhance LLMs’ multilingual capabilities (Ni et al., 2021; Ji et al., 2024). However, these methods typically require substantial investments in time, data, and computational resources. Thus, many researchers perform multilingual alignment to improve LLMs’ multilingual performance by transferring the capabilities from high-resource languages to low-resource languages (Eronen et al., 2023; Zhao et al., 2024c, a; She et al., 2024), which efficiently and effectively improves the model performance in low-resource language scenarios. Furthermore, Zhang et al. (2024) first finds the “Spontaneous Multilingual Alignment” phenomenon in LLMs, which demonstrates that conducting multilingual alignment based on a small number of languages effectively improves the alignment even between English and many languages unseen during alignment.
2.2 Mechanistic Interpretability
@@
In addition to enhancing LLMs’ multilingual performance, research on the underlying mechanisms of multilingual capabilities in LLMs is still ongoing. It is crucial for us to understand and explain the LLMs and related methods explicitly. Typically, the existing approaches primarily perform mechanistic interpretability analyses by observing the internal states of the model (Nostalgebraist, 2020; Zhang et al., 2024; Zhao et al., 2024b; Mousi et al., 2024). Overall, neuron states and latent intermediate logits are both important objects of observation. For latent logits, Wendler et al. (2024) utilizes logit lens (Nostalgebraist, 2020) to directly project the logits in the intermediate layers to the vocabulary space, which reveals the latent participation of English in the intermediate layers. For neuron states, Hu et al. (2024) analyzes the neuron activation overlap to measure the extent of shared neuron activation across different languages.
2.3 Language-Specific Neurons
@@
Many studies have revealed the language-related and language-agnostic components in LLMs. At the layer level, the multilingual processing of LLMs is considered to involve three stages (Zhao et al., 2024b; Wendler et al., 2024): converting multilingual inputs into a shared semantic space, intermediate-layer reasoning, and outputting in the target language. The top and bottom layers of the model handle multilingual processing, while the intermediate layers perform inference in similar patterns across different languages. This demonstrates a distinct division of labor within the model at the layer level regarding language specificity.
@@
Furthermore, many studies investigate the finer-grained methods for language-specific neuron identification (Kojima et al., 2024; Tang et al., 2024). Tang et al. (2024) categorizes activated neurons into language-specific neurons and language-agnostic neurons. They detect language-specific neurons by calculating language activation probability entropy on massive text. However, we find that some neurons are activated by multiple languages (i.e., not language-specific), yet are not universally activated across all languages (i.e., not language-agnostic). Simply categorizing activated neurons into two classes blurs this distinction. Thus, we propose a new method to identify neurons, which categorizes activated neurons into three types: language-specific, language-related, and language-agnostic.
3 Methodology
In this section, we introduce our overall analysis pipeline in our work. First, we review the existing multilingual alignment algorithm and neuron analysis techniques as a preliminary study in §3.1. Then we introduce the multilingual alignment algorithm we utilize in our work in §3.2. Finally, for mechanistic interpretability analysis, we introduce our new method for detecting language-specific, language-related, and language-agnostic neurons (§3.3).
3.1 Preliminary Study
@@
Most LLMs are pretrained mainly on the high-resource language corpus, which leads to LLMs’ unstable and unbalanced performance in multilingual scenarios. As a representative multilingual alignment algorithm, Multilingual-Alignment-as-Preference Optimization (MAPO) (She et al., 2024) effectively and efficiently improves the LLMs’ multilingual performance. Additionally, it is also important for us to understand and analyze the mechanism of LLMs’ multilingual capabilities and multilingual alignment. Moreover, some studies on the identification of the language-specific and language-agnostic neurons in LLMs (Tang et al., 2024; Kojima et al., 2024). It is found that LLMs’ capabilities of processing a particular language mainly come from a small subset of neurons (Tang et al., 2024).
@@
However, there are still many important questions waiting for further investigation. On the one hand, many methods overlook neurons activated by multiple languages yet not language-agnostic, namely language-related neurons that lie between language-specific and language-agnostic categories. On the other hand, research from the perspective of language neurons on the underlying mechanisms of LLMs’ multilingual alignment and multilingual capabilities remains quite limited, which is essential for better understanding and improving the multilingual performance of LLMs.
3.2 Multilingual Alignment
@@
MAPO is a typical multilingual alignment algorithm to align the reasoning capabilities of non-English language responses with those of English, which serves as the pivot language. Specifically, for a given query $X$ in a target (non-English) language and its corresponding English variant $X_{Eng}$, we collect their respective responses $Y$ and $Y_{Eng}$. An off-the-shelf translation model, parameterized by $\theta$, is deployed to estimate the conditional generation probability $P(Y\mid Y_{Eng};\theta)$ by force-decoding $Y$ conditioned on $Y_{Eng}$. A higher conditional probability is interpreted as stronger alignment between the target language response and its English counterpart. This probability is then used as an alignment score, denoted $r_{\theta}(X,Y)$.
@@
This alignment score can be integrated into preference optimization algorithms. For instance, in PPO (Schulman et al., 2017), $r_{\theta}(X,Y)$ can be directly employed as the reward score. In DPO (Rafailov et al., 2023), for each target language, $n$ distinct outputs are generated. Based on the alignment score  these $n$ outputs are used to form $\binom{n}{2}$ preference pairs $(Y_{w},Y_{l})$, where $Y_{w}$ is deemed superior to $Y_{l}$ due to a higher alignment score. The model is then optimized by Eq1.
$L_{DPO}(\pi_{\theta};\pi_{ref})=-E_{(X,Y_{w},Y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(Y_{w}|X)}{\pi_{ref}(Y_{w}|X)}-\beta\log\frac{\pi_{\theta}(Y_{l}|X)}{\pi_{ref}(Y_{l}|X)}\right)\right]$ (1)
3.3 Language Neurons Identification
@@
Following Tang et al. (2024), the neurons in our work are defined as a linear transformation of a single column in a weighted matrix $W$ followed by a non-linear activation, SiLU (Shazeer, 2020). For the $j$-th neuron in the $i$-th layer, its activation probability when processing responses in language $k$ is computed as:
@@
We define language neurons as those exhibiting higher activation probabilities for some languages, while discriminating from others. To select neurons that, despite exhibiting relatively higher entropy, still demonstrate high activation probabilities for some languages, we additionally incorporate the maximum activation probability, as formulated in Eq 3:
@@
where $p_{i,j}^{\prime}$ represents the probability distribution $p_{i,j}$ after normalization and $\lambda$ is a balancing coefficient. Neurons with scores falling in the lowest percentile, specifically, the bottom 5% are selected.
Furthermore, to identify how many languages each selected neuron is related to, we introduce a threshold $\tau$ and compute:
@@
A neuron is considered as a language-specific neuron if $N_{i,j}=1$, and as a language-related neuron if $1<N_{i,j}<l$. Both of them belong to language neuron. In contrast, a neuron is considered language-agnostic neuron if it exhibits high activation probabilities across all $l$ languages.
Finally, given our focus on multilingual reasoning tasks, we select neurons exclusively based on responses from multilingual reasoning datasets, rather than relying on multilingual plain text (Tang et al., 2024).
4 Experiments
4.1 Experimental Setup
Following She et al. (2024), we conduct our experiments and analyses on the mathematical reasoning tasks and different languages. In this section, we introduce our experimental settings in detail.
@@ 4.1 Experimental Setup
We include two different models in our experiments and analyses. Following She et al. (2024), we conduct our experiments on MistralMathOctopus-7B  [^2] https://huggingface.co/kevinpro/MistralMathOctopus-7B and MetaMathOctopus-7B  [^3] https://huggingface.co/kevinpro/MetaMathOctopus-7B. MistralMathOctopus is obtained by fine-tuning MetaMath-Mistral (Yu et al., 2023) with MGSM8KInstruct (Chen et al., 2023). MetaMathOctopus is obtained by fine-tuning MetaMath (Yu et al., 2023) with MGSM8KInstruct. Considering limited computational resources and reproducibility, we directly utilize the publicly released base models. Our analyses are mainly based on MistralMathOctopus in the main text and we report more results in the Appendix.
@@
We conduct experiments on two representative mathematical reasoning benchmarks, MGSM (Shi et al., 2022) and MSVAMP (Chen et al., 2023). MGSM is a widely used benchmark for multilingual mathematical reasoning evaluation. MSVAMP is an out-of-domain test set in contrast to MGSM, which evaluates robustness and generalization (Zhu et al., 2024; She et al., 2024).
@@
Following She et al. (2024), we choose the following 10 different languages for analysis in our work. As a pivot language, English (en) is used as the alignment target. We also choose Chinese (zh), Russian (ru), German (de), French (fr), Spanish (es), Japanese (ja), Swahili (sw), Thai (th) and Bengali (bn) as 9 representative non-English languages.
(a) Base - Language-Specific Neurons
(b) Aligned - Language-Specific Neurons
(c) Base - Language Neurons
(d) Aligned - Language Neurons
Figure 2: PPL changes of MistralMathOctopus on MGSM after deactivating language-specific neurons or language neurons. “Base” indicates the results of the base model. “Aligned” indicates the results of the aligned model. For comparison, the results of Tang et al. (2024) are provided in Appendix D.
@@
Due to limited computational resources, our exploration focuses on the most effective DPO variant of MAPO (She et al., 2024). We select 1, 4, and 8 tasks from the NumGLUE (Mishra et al., 2022), an arithmetic reasoning benchmark, and translate questions into 9 languages, consistent with the MGSM, thereby creating a multilingual seed dataset. To construct preference pairs, we sample responses using the corresponding base models and employ NLLB-200-distilled-600M [^4] https://huggingface.co/facebook/nllb-200-distilled-600M as the translation model to obtain alignment scores. Finally, for each model and each target language (excluding English), we gain 10,000 preference pairs. Training is conducted using LoRA (Hu et al., 2022). During the neuron selection stage, we perform force-decoding on the responses of the MGSM or MSVAMP dataset to obtain the activation probabilities of neurons for each language. Based on empirical results on development sets, we set the balancing coefficient $\lambda=0.04$ and the threshold $\tau=0.5$. Additional implementation details are provided in Appendix B.
4.2 Language Neurons Identification
@@
Based on the neuron identification algorithm introduced in §3.3, we identify the language-specific neurons, language-related neurons, and language-agnostic neurons in the model. To further validate the effectiveness of our algorithm, we follow  Tang et al. (2024) by examining changes in the perplexity of LLMs by deactivating the identified language neurons across different languages. Experiments are conducted on both the base model and the aligned model, with results presented in Figure 2. We report the results of both language-specific neurons and language neurons.
It can be found that whether deactivating language-specific neurons or all language neurons, the results consistently exhibit the same pattern: the diagonal elements in each row show the highest values. Notably, deactivating language neurons leads to a more pronounced effect compared to deactivating only language-specific neurons. These observations support the following findings: (1) Our algorithm effectively identifies language-specific and language-related neurons; (2) For a given language, in addition to its language-specific neurons, there are also a substantial number of shared language-related neurons contributing to its performance; (3) Deactivating all the language-related neurons of one language doesn’t cause significant impacts on the model’s performance in other languages. The above findings confirm the validity of the language neurons identified by our method and further provide insights into the characteristics of language neurons.
4.3 Layer-wise Functionality Analysis
@@
Based on the identified neurons, we perform layer-wise functional analyses of all layers in the LLMs. We begin by analyzing the distributions of different types of neurons in the base model. And we report the results in Figure 3. Through the analysis of the distribution of different types of neurons, we can further divide the LLMs’ internal process for multilingual inference into four parts rather than the three-stage division discussed in some works (Wendler et al., 2024; Zhao et al., 2024b):
@@
    Multilingual Understanding: In the initial layers, the number of language neurons (language-specific and language-related neurons) peaks, while the number of language-agnostic neurons is relatively low. The model maps multilingual inputs into a unified semantic space at this stage.
@@
    Shared Semantic Space Reasoning: In the intermediate layers, the model engages in reasoning within a shared semantic space across different languages. During this stage, language neurons are largely absent, whereas language-agnostic neurons become dominant.

### v2

@@ Abstract
Multilingual Alignment is an effective and representative paradigm to enhance LLMs’ multilingual capabilities, which transfers the capabilities from the high-resource languages to the low-resource languages. Meanwhile, some research on language-specific neurons provides a new perspective to analyze and understand LLMs’ mechanisms. However, we find that there are many neurons that are shared by multiple but not all languages and cannot be correctly classified. In this work, we propose a ternary classification methodology that categorizes neurons into three types, including language-specific neurons, language-related neurons, and general neurons. And we propose a corresponding identification algorithm to distinguish these different types of neurons. Furthermore, based on the distributional characteristics of different types of neurons, we divide the LLMs’ internal process for multilingual inference into four parts: (1) multilingual understanding, (2) shared semantic space reasoning, (3) multilingual output space transformation, and (4) vocabulary space outputting. Additionally, we systematically analyze the models before and after alignment with a focus on different types of neurons. We also analyze the phenomenon of “Spontaneous Multilingual Alignment”. Overall, our work conducts a comprehensive investigation based on different types of neurons, providing empirical results and valuable insights to better understand multilingual alignment and multilingual capabilities of LLMs.
Codehttps://github.com/NJUNLP/Language-Neurons-Alignment
Extended versionhttps://arxiv.org/abs/2505.21505
Introduction
@@
By training on the extensive corpus, large language models (LLMs) demonstrate outstanding language capabilities (Yang et al. 2024; Liu et al. 2024; Grattafiori et al. 2024; Zhang et al. 2025). However, due to the unbalanced pretraining corpus across different languages, LLMs have very uneven performance on high-resource languages and low-resource languages (Huang et al. 2023; Zhu et al. 2023; Zhang et al. 2024; Fan et al. 2025). Therefore, researchers have conducted comprehensive explorations to further enhance the multilingual performance of LLMs. A straightforward approach is increasing the proportion of non-English texts during pretraining (Ni et al. 2021; Yang et al. 2024) or performing continual pretraining with multilingual texts (Liu, Winata, and Fung 2021; Ji et al. 2024). But these approaches often entail high computational costs and substantial amounts of multilingual data.
@@
Figure 1: A neuron’s activation probability across different languages. This neuron that exhibits high activation probabilities across multiple (green) but not all (blue) languages can’t be correctly categorized under the existing classification methodology. The dashed line denotes the threshold that determines whether a neuron exhibits high activation in a given language.
Considering LLMs’ great performance on high-resource languages, multilingual alignment has emerged as a representative paradigm for enhancing multilingual reasoning by transferring knowledge from high-resource to low-resource languages (Zhao et al. 2024a; She et al. 2024). A representative example is MAPO (Multilingual-Alignment-as-Preference Optimization) (She et al. 2024), which improves multilingual alignment by utilizing a well-trained multilingual translation model to compute alignment scores based on the conditional generation probability of translating non-English responses into English.
Many studies conduct systematic mechanism analyses of the multilingual alignment and LLMs’ multilingual capabilities. Zhao et al. (2024b) split the multilingual processing workflow into three parts: multilingual understanding, resolving tasks, and generating outputs in the target language. This three-stage inference workflow clearly demonstrates how LLMs leverage English as a pivot language to handle multilingualism using a unified pattern.
Inspired by the neurobiological underpinnings of human language faculties, Tang et al. (2024) categorize neurons in LLMs into two primary types: language-specific neurons and general neurons. Notably, these language-specific neurons are primarily situated in the model’s top and bottom layers (Tang et al. 2024), which is consistent with the three-stage multilingual workflow of Zhao et al. (2024b).
However, we identify a key limitation in the existing neuron classification methodology: the identification of language-specific neurons focuses on the processing a particular language and thus neglects the inter-language alignment; while general neurons encode universal knowledge which is independent of a specific language. This leads to a critical question: what is the mechanism of sharing neurons between some languages? Furthermore, it is not well understood how multilingual alignment enhances the LLMs’ multilingual reasoning capabilities from the perspective of language neurons.
In this work, we comprehensively investigate the multilingual alignment of LLMs with MAPO (She et al. 2024) as a representative multilingual alignment algorithm. As shown in Figure 1, we observed the existence of neurons that exhibit high activation probability on multiple but not all languages. These neurons are neither language-specific nor general for all languages. Hence, we refer to them as language-related neurons.
Moreover, we redefine language-specific neurons to restrict each neuron to be activated for only one language. The general neurons are defined as the neurons that are effective for all languages. To facilitate subsequent analysis, we propose a corresponding identification algorithm for distinguishing these different types of neurons.
@@
Then we analyze the models before and after alignment, focusing on the changes in different types of neurons. Based on the distributional characteristics of these neurons, we divide LLMs’ internal process for multilingual inference into four parts, with different parts exhibit distinct dependencies on different types of neurons. We demonstrate that multilingual alignment significantly enhances the activation of the corresponding types of neurons across the relevant layers. Additionally, we analyze the “spontaneous multilingual alignment” (Zhang et al. 2024) phenomenon in LLMs, providing insights into the roles of general neurons and language-related neurons shared across languages. For further analysis, we also provide observations about the uniqueness of English and the neuron distributions. Overall, based on different types of neurons, we present empirical results and valuable insights that contribute to a deeper understanding of multilingual alignment and the multilingual reasoning capabilities of LLMs.
Related Work
Multilingual Alignment
@@
Conducting pretraining or continual pretraining on the multilingual corpus is a straightforward and effective method to enhance LLMs’ multilingual capabilities (Ni et al. 2021; Ji et al. 2024). However, these methods typically require substantial investments in time, data, and computational resources. Thus, many researchers perform multilingual alignment to improve LLMs’ multilingual performance by transferring the capabilities from high-resource languages to low-resource languages (Eronen, Ptaszynski, and Masui 2023; Zhao et al. 2024c, a; She et al. 2024), which efficiently and effectively improves the model performance in low-resource language scenarios. Furthermore, Zhang et al. (2024) first finds the “Spontaneous Multilingual Alignment” phenomenon in LLMs, which demonstrates that conducting multilingual alignment based on a small number of languages effectively improves the alignment even between English and many languages unseen during alignment.
Mechanistic Interpretability
@@
In addition to enhancing LLMs’ multilingual performance, research on the underlying mechanisms of multilingual capabilities in LLMs is still ongoing. It is crucial for us to understand and explain the LLMs and related methods explicitly. Typically, the existing approaches primarily perform mechanistic interpretability analyses by observing the internal states of the model (Nostalgebraist 2020; Zhang et al. 2024; Zhao et al. 2024b; Mousi et al. 2024). Overall, neuron states and latent intermediate logits are both important objects of observation. For latent logits, Wendler et al. (2024) utilizes logit lens (Nostalgebraist 2020) to directly project the logits in the intermediate layers to the vocabulary space, which reveals the latent participation of English in the intermediate layers. For neuron states, Hu et al. (2024) analyzes the neuron activation overlap to measure the extent of shared neuron activation across different languages.
Language-Specific Neurons
@@
Many studies have revealed the language-related and universal components in LLMs. At the layer level, the multilingual processing of LLMs is considered to involve three stages (Zhao et al. 2024b; Wendler et al. 2024): converting multilingual inputs into a shared semantic space, intermediate-layer reasoning, and outputting in the target language. The top and bottom layers of the model handle multilingual processing, while the intermediate layers perform inference in similar patterns across different languages. This demonstrates a distinct division of labor within the model at the layer level regarding language specificity.
@@
Furthermore, many studies investigate the finer-grained methods for language-specific neuron identification (Kojima et al. 2024; Tang et al. 2024; Tan, Wu, and Monz 2024). Tang et al. (2024) categorizes activated neurons into language-specific neurons and general neurons. They detect language-specific neurons by calculating language activation probability entropy on massive text. However, we find that some neurons are activated by multiple languages (i.e., not language-specific), yet are not universally activated across all languages (i.e., not general). Simply categorizing activated neurons into two classes blurs this distinction. Thus, we propose a new method to identify neurons, which categorizes activated neurons into three types: language-specific neurons, language-related neurons, and general neurons.
Methodology
Preliminary Study
@@
Most LLMs are pretrained mainly on the high-resource language corpus, which leads to LLMs’ unstable and unbalanced performance in multilingual scenarios. As a representative multilingual alignment algorithm, Multilingual-Alignment-as-Preference Optimization (MAPO) (She et al. 2024) effectively and efficiently improves the LLMs’ multilingual performance. Additionally, it is also important for us to understand and analyze the mechanism of LLMs’ multilingual capabilities and multilingual alignment. Moreover, some studies on the identification of the language-specific and general neurons in LLMs (Tang et al. 2024; Kojima et al. 2024). It is found that LLMs’ capabilities of processing a particular language mainly come from a small subset of neurons (Tang et al. 2024).
@@
However, there are still many important questions waiting for further investigation. On the one hand, many methods overlook neurons activated by multiple languages yet not general, namely language-related neurons lie between language-specific and general categories. On the other hand, research from the perspective of language neurons on the underlying mechanisms of LLMs’ multilingual alignment and multilingual reasoning capabilities remains quite limited, which is essential for better understanding and improving the multilingual performance of LLMs.
Multilingual Alignment
@@
MAPO is a typical multilingual alignment algorithm to align the reasoning capabilities of non-English language responses with those of English, which serves as the pivot language. Specifically, for a given query $X$ in a target (non-English) language and its corresponding English variant $X_{En}$, we collect their respective responses $Y$ and $Y_{En}$. An off-the-shelf translation model, parameterized by $\theta$, is deployed to estimate the conditional generation probability $P(Y\mid Y_{En};\theta)$ by force-decoding $Y$ conditioned on $Y_{En}$. A higher conditional probability is interpreted as stronger alignment between the target language response and its English counterpart. This probability is then used as an alignment score, denoted $r_{\theta}(X,Y)$.
@@
This alignment score can be integrated into preference optimization algorithms. In PPO (Schulman et al. 2017), $r_{\theta}(X,Y)$ can be directly employed as the reward score. In DPO (Rafailov et al. 2023), for each target language, $n$ distinct outputs are generated. Based on the alignment score, these $n$ outputs are used to form $\binom{n}{2}$ preference pairs $(Y_{w},Y_{l})$, where $Y_{w}$ is deemed superior to $Y_{l}$ due to a higher alignment score. The model is optimized by Eq. 1 to Eq. 3:
$A$ $=\beta\log\frac{\pi_{\theta}(Y_{w}|X)}{\pi_{ref}(Y_{w}|X)}$ (1)
$B$ $=\beta\log\frac{\pi_{\theta}(Y_{l}|X)}{\pi_{ref}(Y_{l}|X)}$ (2)
$L_{DPO}(\pi_{\theta};\pi_{ref})$ $=-E_{(X,Y_{w},Y_{l})\sim D}\left[\log\sigma\left(A-B\right)\right]$ (3)
(a) Language-specific Neurons
(b) Language-specific & Language-related Neurons
Figure 2: PPL changes of MistralMathOctopus on MGSM after deactivating language-specific neurons or language-specific & language-related neurons. For comparison, the results of Tang et al. (2024) are provided in appendix.
Language Neurons Identification
@@
Following Tang et al. (2024), the neurons in our work are defined as a linear transformation of a single column in a weighted matrix $W$ followed by a non-linear activation, SiLU (Shazeer 2020). For the $j$-th neuron in the $i$-th layer, its activation probability when processing responses in language $k$ is computed as:
@@
We define language-related neurons as those neurons exhibiting high activation probabilities for multiple but not all languages. However, methods that only focus on language specificity are defective in detecting the above neurons. To better detect the different types of neurons, we trade off two intrinsic properties of each neuron: (1) Language-specificity, which is quantified by the entropy of its activation probability distribution across languages; (2) Effectiveness, which measures the extent to which the neuron participates when the model solves tasks. And it is quantified by the neuron’s maximum activation probability across different languages in our work. By doing so, we are better able to simultaneously identify neurons that are relatively more specific and relatively more generalized. They are combined into a unified metric as formulated in Eq 5:
@@
where $p_{i,j}^{\prime}$ represents the probability distribution $p_{i,j}$ after normalization and $\lambda$ is a balancing coefficient. Specifically, we automatically determine a value of $\lambda$ such that general neurons are not identified as language-specific or language-related neurons shared by all $l$ languages. Following Tang et al. (2024), neurons with scores falling in the lowest $1\%$ are selected.
Furthermore, to identify how many languages each selected neuron is related to, we introduce a threshold $\tau$:
@@
Following Tang et al. (2024), the threshold $\tau$ is set to the top 5% of all activation probabilities. A neuron is considered as a language-specific neuron if $N_{i,j}=1$, and as a language-related neuron if $1<N_{i,j}<l$. Meanwhile, a neuron is considered general neuron if it exhibits high activation probabilities across all $l$ languages.
Finally, given our focus on multilingual reasoning tasks, we select neurons exclusively based on responses from multilingual reasoning datasets, rather than relying on multilingual plain text (Tang et al. 2024).
Experiments
Experimental Setup
Following She et al. (2024), we conduct our experiments and analyzes on the mathematical reasoning tasks in different languages. In this section, we introduce our experimental settings in detail.
@@
We include two different models in our experiments and analyses. Following She et al. (2024), we conduct our experiments on MistralMathOctopus-7B  [^1] https://huggingface.co/kevinpro/MistralMathOctopus-7B and MetaMathOctopus-7B  [^2] https://huggingface.co/kevinpro/MetaMathOctopus-7B. MistralMathOctopus is obtained by fine-tuning MetaMath-Mistral (Yu et al. 2023) with MGSM8KInstruct (Chen et al. 2023). MetaMathOctopus is obtained by fine-tuning MetaMath (Yu et al. 2023) with MGSM8KInstruct. Considering limited computational resources and reproducibility, we directly utilize the publicly released base models. Our analyses are mainly based on MistralMathOctopus in the main text and we report more results in the appendix.
@@
We conduct experiments on two representative mathematical reasoning benchmarks, MGSM (Shi et al. 2022) and MSVAMP (Chen et al. 2023). MGSM is a widely used benchmark for multilingual mathematical reasoning evaluation. MSVAMP is an out-of-domain test set in contrast to MGSM, which evaluates robustness and generalization (Zhu et al. 2024; She et al. 2024).
@@
Following She et al. (2024), we choose the following 10 different languages for analysis. As a pivot language, English (en) is used as the alignment target. We also choose Chinese (zh), Russian (ru), German (de), French (fr), Spanish (es), Japanese (ja), Swahili (sw), Thai (th) and Bengali (bn) as 9 representative non-English languages.
@@
Due to limited computational resources, our exploration focuses on the most effective DPO variant of MAPO (She et al. 2024). We select 1, 4, and 8 tasks from the NumGLUE (Mishra et al. 2022), an arithmetic reasoning benchmark, and translate questions into 9 languages, consistent with the MGSM, thereby creating a multilingual seed dataset. To construct preference pairs, we sample responses using the corresponding base models and employ NLLB-200-distilled-600M [^3] https://huggingface.co/facebook/nllb-200-distilled-600M as the translation model to obtain alignment scores. Finally, for each model and each target language (excluding English), we gain 10,000 preference pairs. Training is conducted using LoRA (Hu et al. 2022). During the neuron selection stage, we perform force-decoding on the responses of the MGSM or MSVAMP dataset to obtain the activation probabilities of neurons for each language. Additional implementation details are provided in appendix.
Language Neurons Identification
@@
Based on the neuron identification algorithm, we identify the language-specific neurons, language-related neurons, and general neurons in the model. To further validate the effectiveness of our algorithm, we follow  Tang et al. (2024) by examining changes in the perplexity (PPL) scores of LLMs by deactivating the identified neurons across different languages. Experiments are conducted on the base model, with results presented in Figure 2. We report the results of two settings, deactivating language-specific neurons and deactivating language-specific & language-related neurons. The higher PPL change (darker cells) indicates the stronger reliance on the neurons be deactivated.
It can be found that the performance for each language mostly relies on both its language-specific neurons and language-related neurons rather than other neurons, as the diagonal elements in each row show the highest changes in PPL. However, different languages vary in the extent to which they rely on their corresponding language-specific and language-related neurons, reflecting differences in cross-lingual alignment across different languages.
Notably, deactivating both language-specific and language-related neurons leads to a more pronounced effect compared to deactivating only language-specific neurons. Suggesting that for a given language, in addition to its language-specific neurons, shared language-related neurons also significantly contribute to its performance.
Also, deactivating all the language-related neurons of one language doesn’t cause significant impacts on the model’s performance in other languages, which indicates that the language-related neurons for a specific language are relatively evenly dispersed across multiple other languages. The above findings confirm the validity of the language neurons identified by our method and further provide insights into the characteristics of language neurons.
Layer-wise Functionality Analysis
@@
Based on the identified neurons, we perform layer-wise functional analyses of all layers in the LLMs. We begin by analyzing the distributions of different types of neurons in the base model. And we report the results in Figure 3.
Some works have divided the LLM’s multilingual inference process into three stages (Wendler et al. 2024; Zhao et al. 2024b). However, through the analysis of the distribution of different types of neurons, we find that although the combined function of the last stage is still to produce language-specific symbols, the behavior of language-specific neurons first reaches its peaks and then declines, which differs substantially from the other two types. So we further divide the final stage into two parts, leading to a four-stage interpretation of the LLMs’ internal process for multilingual inference:
@@
    Multilingual Understanding: In the initial layers, the number of both language-specific and language-related neurons peaks, while the number of general neurons is relatively low. The model maps multilingual inputs into a unified semantic space at this stage.
@@
    Shared Semantic Space Reasoning: In the intermediate layers, the model engages in reasoning within a shared semantic space across different languages. During this stage, both language-specific and language-related neurons are largely absent, whereas general neurons become dominant.